OpenStudy (anonymous):

2^(a+1)=9 2^(b-1)=27 write a as a product of B? answer = (2b-5)/3 How can i solve this? Thanks

OpenStudy (anonymous):

a+1=2log(3) b-1=3log (3) (2b-2)/3=2log3=a+1 a=(2b-5)/3 Log is taken with the base 2